# Density-based User Representation using Gaussian Process Regression for Multi-interest Personalized Retrieval

### v4

@@ 1 Introduction
With the proliferation of online platforms, users have ready access to content, products and services drawn from a vast corpus of candidates. Personalized recommender systems (RSs) play a vital role in reducing information overload and helping users navigate this space. It is widely recognized that users rarely have a single intent or interest when interacting with an RS (Weston et al., 2013; Pal et al., 2020; Cen et al., 2020). To enhance personalization, recent work focuses on discovering a user’s multiple interests and recommending items that attempt to span their interests (Cen et al., 2020; Tan et al., 2021; Li et al., 2019). However, this is challenging for two reasons. First, user interests are diverse and dynamic: diversity makes it hard to detect all interests, while their dynamic nature renders determining which user interest is active at any given time quite difficult. Second, it is hard to retrieve items related to niche interests due to the popularity bias (Chen et al., 2020).
User representation is a fundamental design choice in any RS. A key component in capturing multiple interests is a representation that naturally encodes a user’s diverse preferences. The most widely used strategy for user modeling is the single-point user representation (SUR), which uses a single point in an item embedding space to represent the user. The user’s affinity for an item is obtained using some distance measure (e.g., inner product, cosine similarity) with the point representing the item. However, SUR can limit the accuracy and diversity of item retrieval (Zhang et al., 2023); hence, most RSs generally use high-dimensional embedding vectors (with high computation cost).
@@
To address the limitations of SUR, MaxMF (Weston et al., 2013) adopts a multi-point user representation (MUR), where each user is represented using $K$ points in the embedding space, each reflecting a different “interest”. MaxMF uses a constant, uniform $K$ across all users (e.g., $K=4$), which is somewhat ad hoc and very restrictive. Subsequent research uses other heuristics (Liu et al., 2019; Cen et al., 2020; Wang et al., 2020a, b; Tan et al., 2021; Chen et al., 2021; Liu et al., 2022) or clustering algorithms (Li et al., 2019; Pal et al., 2020) to determine the number of interests per user. However, these all require the manual choice of $K$ or a specific clustering threshold, limiting the adaptability of MUR methods, since interests generally have high variability across users. Moreover, uncertainty regarding a user’s interests is not well-modeled by these methods, diminishing their ability to perform effective online exploration.
@@
To address limitations of SUR and MUR point-based representations, we propose a user representation that emphasizes (i) adaptability, adapting to different interest patterns; (ii) uncertainty-awareness, modeling uncertainty in assessing user interests; and (iii) efficiency, avoiding high-dimensional embeddings. Specifically, we use a density-based representation, where the user’s preferences are encoded using a function over the item embedding space. Under this representation, the relevance score for user-item pairs should be higher in regions of embedding space where a user demonstrated more interest in the corresponding items, and lower in regions where users have shown limited interest. We propose the density-based user representation (DUR), a novel user modeling method which exploits Gaussian process regression (GPR) (Rasmussen, 2004; Rasmussen & Williams, 2005), a well-studied Bayesian approach to non-parametric regression using Gaussian processes (GPs) to extrapolate from training to test data. GPR has been applied across a wide range of domains, though it has been under-explored in user modeling. Given a sequence of user interactions, GPR predicts the probability of a user’s interest in any item using its posterior estimates. This allows us to maintain a unique GP regressor for each user, effectively capturing their evolving preferences and assessing uncertainty. Top-$N$ item retrieval is performed using bandit algorithms, such as UCB (Auer, 2003) or Thompson sampling (Russo et al., 2018), based on GPR posterior estimates.
@@
To illustrate, consider Fig. 1, which shows the prediction score (i.e., the inner-product between user and item embeddings) between a user and all movies from the MovieLens 1M dataset (Harper & Konstan, 2015) (reduced to 2D for visualization purposes). We examine 20 movies from the recent history of a particular user, shown as triangles ($\blacktriangle$). We see that these movies lie in several different regions of the embedding space. However, when we fit either SUR or MUR ($K=4$) models, they fail to capture the user’s multiple interests and instead assign high scores only to movies from a single region (Fig. 1, top row). By contrast, we see in Fig. 1 (bottom left) that GPR fits the data well, assigning high values to all regions associated with the user’s recent watches (interests). Fig. 1 (bottom right) shows that our approach can also capture uncertainty in our estimates of a user’s interests, assigning high uncertainty to regions in embedding space with fewer samples.
@@ 2 Related Work
Learning high-quality user representations is central to good RS performance. The single-point user representation (SUR) is the dominant approach, where a user is captured by a single point in some embedding space (Ricci et al., 2011; Rendle et al., 2009), for example, as employed by classical (Koren, 2008; Rendle, 2010; Chen et al., 2018) and neural (He et al., 2017) collaborative filtering (CF) methods. While effective and widely used, SUR cannot reliably capture a user’s multiple interests.
@@
To address this limitation, the multi-point user representation (MUR) (Weston et al., 2013) has been proposed, where a user is represented by multiple points in embedding space, each corresponding to a different “primary” interest. Selecting a suitable number of points $K$ is critical in MUR. Existing algorithms largely use heuristic methods, e.g., choosing a global constant $K$ for all users (Weston et al., 2013; Liu et al., 2019; Wang et al., 2020a; Cen et al., 2020; Wang et al., 2020b; Tan et al., 2021; Chen et al., 2021; Liu et al., 2022). Other methods personalize $K$ by setting it to the logarithm of the number of items with which a user has interacted (Li et al., 2019). More recently, Ward clustering of a user’s past items has been proposed, with a user’s $K$ determined by the number of such clusters (Pal et al., 2020). This too requires manual tuning of clustering thresholds.
@@
At inference time MUR is similar to SUR, computing the inner-product of the user embedding(s) and item embedding. Some methods compute $K$ inner products, one per interest, and use the maximum as the recommendation (and the predicted score for that item) (Weston et al., 2013). Others first retrieve the top-$N$ items for each interest ($N\times K$ items), then recommend the top-$N$ items globally (Cen et al., 2020; Pal et al., 2020). None of these methods capture model uncertainty w.r.t. a user’s interests, hence they lack the ability to balance exploration and exploitation in online recommendation in a principled way (Chen, 2021).
@@
Our density-based user representation, and our proposed GPR4DUR, differs from prior work w.r.t. both problem formulation and methodology. Most prior MUR methods focus on next-item prediction (Cen et al., 2020; Liu et al., 2022), implicitly assuming a single-stage RS, where the trained model is the main recommendation engine. However, many practical RSs consist of two stages: candidate selection (or retrieval) followed by ranking (Covington et al., 2016b; Eksombatchai et al., 2018). This naturally raises the question: are the selected candidates diverse enough to cover a user’s interests or intents? This is especially relevant when a user’s dominant interest at the time of recommendation is difficult to discern with high probability; hence, it is important that the ranker have access to a diverse set of candidates that cover the user’s range of potential currently active interests. In this paper, we focus on this retrieval task. As for methodology, almost all prior work uses SUR or MUR point-based representations,  (Weston et al., 2013; Pal et al., 2020; Cen et al., 2020; Li et al., 2019)—these fail to satisfy all the desiderata outlined in Sec. 1. Instead, we propose DUR), a novel method satisfying these criteria, and, to the best of our knowledge, the first to adopt GPR for user modeling in multi-interest recommendation/retrieval.
@@
A related non-parametric recommendation approach is $k$-nearest-neighbors (kNN), where users with similar preferences are identified (e.g., in user embedding space), and their ratings generate recommendations (see, e.g., Grcar (2004)). Our approach differs by not using user embeddings, but fitting a GPR model directly to item embeddings (see Sec. 4.4), allowing for uncertainty in the user model.
@@ 3 Formulation and Preliminaries
In this section, we outline our notation and multi-interest retrieval problem formulation, and provide some background on GPR, which lies at the core of our DUR model.
Notation. We consider a scenario where each recommended item is associated with category information (e.g., genre for movies). Denote the set of all users, items, and categories by $\mathcal{U}$, $\mathcal{V}$, and $\mathcal{C}$, respectively. For each $u\in\mathcal{U}$, whose interaction history has length $l_{u}$, we partition the sequence of items $\mathcal{V}_{u}$ in $u$’s history into two disjoint lists based on the interaction timestamp (which are monotonic increasing): (i) the history set $\mathcal{V}^{\text{h}}_{u}=[v_{u,1},v_{u,2},...,v_{u,\ell_{u}}]$ serves as the model input; and (ii) the holdout set $\mathcal{V}^{\text{d}}_{u}=[v_{u,{\ell_{u}+1}},v_{u,{\ell_{u}+2}},...,v_{u,l_{u}}]$ is used for evaluation. We define $u$’s interests $\mathcal{C}(\mathcal{V}_{u})$ to be the set of categories associated with all items in $u$’s history. Our notation is summarized in Table 4 in the appendix.
@@
We note that our problem is related to sequential recommendation, where the input is a sequence of interacted items sorted by the timestamp per user, and the goal is to predict the next item with which the user will interact. However, here we focus on retrieving a set of items that cover a user’s diverse interests. We defer the task of generating a precise recommendation list to a downstream ranking model.
GPR. The core of our model is Gaussian process regression (GPR). A Gaussian Process (GP) is a non-parametric model for regression and classification. GPR models a distribution over functions, providing not only function value estimates for any input, but also uncertainty quantification via predictive variance, enhancing robust decision-making and optimization (Williams & Rasmussen, 1995).
The key components of the GP are the mean function and the covariance (kernel) function, which capture the prior assumptions about the function’s behavior and the relationships between input points, respectively (Rasmussen & Williams, 2005). Let $\mathbf{X}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{n}\}\in\mathbb{R}^{n\times d}$ be a set of input points and $\mathbf{y}=\{y_{1},\dots,y_{n}\}\in\mathbb{R}^{n}$ be the corresponding output values. A GP is defined as:
$f\sim\mathcal{GP}(\mu,k),$ (1)
where $\mu(\mathbf{x})$ is the mean and $k(\mathbf{x},\mathbf{x^{\prime}})$ the covariance function. The joint distribution of the observation and the output at a new point is derived using a Gaussian distribution, as illustrated in the Appendix.
@@ 4 Methodology
We now outline our density-based user representation using GPR and its application to multi-interest retrieval.
@@ 4.1 GPR for Density-based User Representation
We use GPR to construct a novel density-based user representation (DUR) for multi-interest modeling in RSs. Our key insight involves using GPR to learn a DUR, using a user’s interaction history, that naturally embodies their diverse interest patterns. For any user $u$, let $\mathbf{V}_{u}=[\mathbf{v}_{u,1},\mathbf{v}_{u,2},...,\mathbf{v}_{u,l_{u}}]\in\mathbb{R}^{l_{u}\times d}$ be the embeddings of all items in their history. This is derived from their interaction list $\mathcal{V}_{u}$ and an item embedding matrix $\mathbf{V}$ (we describe how to obtain $\mathbf{V}$ in Sec. 4.4). Let $\mathbf{o}_{u}=[o_{u,v}|v\in\mathcal{V}_{u}]\in\mathbb{R}^{l_{u}}$ be the vector of $u$’s observed interactions with items in $\mathcal{V}_{u}$. We employ a GP to model $u$’s interests given the input points $\mathbf{V}_{u}$ and corresponding observations $\mathbf{o}_{u}$ (analogous to Eq. 1), $g_{u}\sim\mathcal{GP}(\mu_{u},k_{u})$, where, $g_{u}$, $\mu_{u}$ and $k_{u}$ are $u$’s personalized GP regressor, mean, and kernel function, respectively. The joint distribution of observation $\mathbf{o}_{u}$ and the predicted observation of a novel item $v_{*}$ is:
@@
$g_{u}(\mathbf{v})|\mathbf{o}_{u}\sim\mathcal{N}(\bar{g}_{u},\mathrm{cov}(g_{u})),$ (3)
@@
$\mathrm{cov}(g_{u})=$ $k(\mathbf{v},\!\mathbf{v})\!+\!\sigma_{*}^{2}\!$
$\,-\!\mathbf{k}(\mathbf{v}_{*},\!\mathbf{V}_{u})[\mathbf{K}(\mathbf{V}_{u},\!\mathbf{V}_{u})\!+\!\sigma^{2}\!\mathbf{I}]^{-1}\mathbf{k}(\mathbf{V}_{u},\!\mathbf{v})^{T}\!.$ (5)
@@ 4.2 Retrieval List Generation
After obtaining a DUR $g_{u}$ for $u\in\mathcal{U}$ using GPR, we generate the retrieval list using the posterior $g_{u}(\mathbf{v})$ over all unobserved items. The top N items with the highest values form our retrieval list. We consider two methods for selection: (i) Thompson sampling (TS), a probabilistic method that selects items based on posterior sampling (Russo et al., 2018) and (ii) Upper Confidence Bound (UCB), a deterministic method that selects items based on their estimated rewards and uncertainties (Auer, 2003; Auer et al., 2002).
@@ 4.3 GPR Parameter Tuning
The free parameters in our model are the kernel function $\mathbf{K}$ and variance $\sigma^{2}$ in Eq. 2. We treat these as hyperparameters of GPR which are optimized using evaluation on a separate holdout set. Specifically, to generate $\mathcal{R}_{u}$ for user $u$, we fit the GPR model not to the complete interaction history $\mathcal{V}_{u}$, but to the reduced history $\mathcal{V}^{\text{h}}_{u}$, using the item embeddings and observed ratings. We assess retrieval performance using specific metrics (see Sec. 5.3) on both $\mathcal{R}_{u}$ and the holdout set $\mathcal{V}^{\text{d}}_{u}$ We tune GPR parameters using these criteria (we detail the adjustable parameters in Sec. 5.2).

### v5

@@ 1 Introduction
With the proliferation of online platforms, users have ready access to content, products and services drawn from a vast corpus of candidates. Personalized recommender systems (RSs) play a vital role in reducing information overload and helping users navigate this space. It is widely recognized that users rarely have a single intent or interest when interacting with an RS [1, 2, 3]. To enhance personalization, recent work focuses on discovering a user’s multiple interests and recommending items that attempt to span their interests [3, 4, 5]. However, this is challenging for two reasons. First, user interests are diverse and dynamic: diversity makes it hard to detect all interests, while their dynamic nature renders determining which user interest is active at any given time quite difficult. Second, it is hard to retrieve items related to niche interests due to the popularity bias [6].
User representation is a fundamental design choice in any RS. The most widely used strategy for user modeling is the single-point user representation (SUR), which uses a single point in an item embedding space to represent the user. The user’s affinity for an item is obtained using some distance measure (e.g., inner product, cosine similarity) with the point representing the item. However, SUR can limit the accuracy and diversity of item retrieval [7]; hence, most RSs generally use high-dimensional embedding vectors (with high computation cost).
@@
To address the limitations of SUR, MaxMF [1] adopts a multi-point user representation (MUR), where each user is represented using $K$ points in the embedding space, each reflecting a different “interest”. MaxMF uses a constant, uniform $K$ across all users (e.g., $K=4$), which is somewhat ad hoc and restrictive. Subsequent research uses other heuristics [3, 8, 9, 4, 10, 11] or clustering algorithms [5, 2] to determine the number of interests per user. However, these all require the manual choice of $K$ or a specific clustering threshold, limiting the adaptability of MUR methods, since interests generally have high variability across users. Moreover, uncertainty regarding a user’s interests is not well-modeled by these methods, diminishing their ability to perform effective online exploration.
@@
To address limitations of SUR and MUR point-based representations, we propose a user representation that emphasizes (i) adaptability, adapting to different interest patterns; (ii) uncertainty-awareness, modeling uncertainty in assessing user interests; and (iii) efficiency, avoiding high-dimensional embeddings. Specifically, we use a density-based representation, where the user’s preferences are encoded using a function over the item embedding space. Under this representation, the relevance score for user-item pairs should be higher in regions of embedding space where a user demonstrated more interest in the corresponding items, and lower in regions where users have shown limited interest. We propose the density-based user representation (DUR), a novel user modeling method which exploits Gaussian process regression (GPR) [12, 13], a well-studied Bayesian approach to non-parametric regression using Gaussian processes (GPs) to extrapolate from training to test data. GPR has been applied across a wide range of domains, though it has been under-explored in user modeling. Given a sequence of user interactions, GPR predicts the probability of a user’s interest in any item using its posterior estimates. This allows us to maintain a unique GP regressor for each user, effectively capturing their evolving preferences and assessing uncertainty. Top-$N$ item retrieval is performed using bandit algorithms, such as UCB [14] or Thompson sampling [15], based on GPR posterior estimates.
@@
To illustrate, consider Fig. 1, which shows the prediction score (i.e., the inner-product between user and item embeddings) between a user and all movies from the MovieLens 1M dataset [16] (reduced to 2D for visualization purposes). We examine 20 movies from the recent history of a particular user, shown as triangles ($\blacktriangle$). We see that these movies lie in several different regions of the embedding space. However, when we fit either SUR or MUR ($K=4$) models, they fail to capture the user’s multiple interests and instead assign high scores only to movies from a single region (Fig. 1, top row). By contrast, we see in Fig. 1 (bottom left) that GPR fits the data well, assigning high values to all regions associated with the user’s recent watches (interests). Fig. 1 (bottom right) shows that our approach can also capture uncertainty in our estimates of a user’s interests, assigning high uncertainty to regions in embedding space with fewer samples.
@@ 2 Related Work
Learning high-quality user representations is central to good RS performance. The single-point user representation (SUR) is the dominant approach, where a user is captured by a single point in some embedding space [17, 18], for example, as employed by classical [19, 20, 21] and neural [22] collaborative filtering methods. While effective and widely used, SUR cannot reliably capture a user’s multiple interests. To address this limitation, the multi-point user representation (MUR) [1] has been proposed, where a user is represented by multiple points in embedding space, each corresponding to a different “primary” interest. Selecting a suitable number of points $K$ is critical in MUR. Existing algorithms largely use heuristic methods, e.g., choosing a global constant $K$ for all users [1, 8, 3, 9, 4, 10, 11]. Other methods personalize $K$ by setting it to the logarithm of the number of items with which a user has interacted [5]. More recently, Ward clustering of a user’s past items has been proposed, with a user’s $K$ determined by the number of such clusters [2]. This too requires manual tuning of clustering thresholds.
@@
At inference time MUR is similar to SUR, computing the inner-product of the user embedding(s) and item embedding. Some methods compute $K$ inner products, one per interest, and use the maximum as the recommendation (and the predicted score for that item) [1]. Others first retrieve the top-$N$ items for each interest ($N\times K$ items), then recommend the top-$N$ items globally [3, 2]. None of these methods capture model uncertainty w.r.t. a user’s interests, hence they lack the ability to balance exploration and exploitation in online recommendation in a principled way [23].
@@
Our density-based user representation, and our proposed GPR4DUR, differs from prior work w.r.t. both problem formulation and methodology. Most prior MUR methods focus on next-item prediction [3, 11], implicitly assuming a single-stage RS, where the trained model is the main recommendation engine. However, many practical RSs consist of two stages: candidate selection (or retrieval) followed by ranking [24, 25]. This naturally raises the question: are the selected candidates diverse enough to cover a user’s interests or intents? This is especially relevant when a user’s dominant interest at the time of recommendation is difficult to discern with high probability; hence, it is important that the ranker have access to a diverse set of candidates that cover the user’s range of potential currently active interests. In this paper, we focus on this retrieval task. As for methodology, almost all prior work uses SUR or MUR point-based representations,  [1, 2, 3, 5]—these fail to satisfy all the desiderata outlined in Sec. 1. Instead, we propose DUR, a novel method satisfying these criteria, and, to the best of our knowledge, the first to adopt GPR for user modeling in multi-interest recommendation/retrieval. We frame our solution as a candidate generator to be used in the retrieval phase of a RS. Other candidate generators with different objectives can be used in parallel to ours.
@@
A related non-parametric recommendation approach is $k$-nearest-neighbors (kNN), where users with similar preferences are identified (e.g., in user embedding space), and their ratings generate recommendations (see, e.g., [26]). Our approach differs by not using user embeddings, but fitting a GPR model directly to item embeddings (see Sec. 4.4), allowing for uncertainty in the user model.
@@ 3 Formulation and Preliminaries
In this section, we outline our notation and multi-interest retrieval problem formulation, and provide some background on GPR, which lies at the core of our DUR method.
Notation. We consider a scenario where each item is associated with category information (e.g., genre for movies). Denote the set of all users, items, and categories by $\mathcal{U}$, $\mathcal{V}$, and $\mathcal{C}$, respectively. For each $u\in\mathcal{U}$, whose interaction history has length $l_{u}$, we partition the sequence of items $\mathcal{V}_{u}$ in $u$’s history into two disjoint lists based on the interaction timestamp (which are monotonic increasing): (i) the history set $\mathcal{V}^{\text{h}}_{u}=[v_{u,1},v_{u,2},...,v_{u,\ell_{u}}]$ serves as the model input; and (ii) the holdout set $\mathcal{V}^{\text{d}}_{u}=[v_{u,{\ell_{u}+1}},v_{u,{\ell_{u}+2}},...,v_{u,l_{u}}]$ is used for evaluation. We define $u$’s interests $\mathcal{C}(\mathcal{V}_{u})$ to be the set of categories associated with all items in $u$’s history. Our notation is summarized in Appendix A.1.
@@
Our problem is related to sequential recommendation, where the input is a sequence of interacted items sorted by the timestamp per user, and the goal is to predict the next item with which the user will interact. However, here we focus on retrieving a set of items that cover a user’s diverse interests. We defer the task of generating a precise recommendation list to a downstream ranking model.
GPR. The core of our model is Gaussian process regression (GPR). A Gaussian Process (GP) is a non-parametric model for regression and classification. GPR models a distribution over functions, providing not only function value estimates for any input, but also uncertainty quantification via predictive variance, enhancing robust decision-making and optimization [27].
The key components of the GP are the mean function and the covariance (kernel) function, which capture the prior assumptions about the function’s behavior and the relationships between input points, respectively [13]. Let $\mathbf{X}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{n}\}\in\mathbb{R}^{n\times d}$ be a set of input points and $\mathbf{y}=\{y_{1},\dots,y_{n}\}\in\mathbb{R}^{n}$ be the corresponding output values. A GP is defined as: $f\sim\mathcal{GP}(\mu,k)$, where $\mu(\mathbf{x})$ is the mean and $k(\mathbf{x},\mathbf{x^{\prime}})$ the covariance function. The joint distribution of the observation and the output at a new point is derived using a Gaussian distribution, as illustrated in Appendix A.1.
@@ 4 Methodology
We outline density-based user representation using GPR and its application to multi-interest retrieval.
@@ 4.1 GPR for Density-based User Representation
We use GPR to construct a novel density-based user representation (DUR) for multi-interest modeling in RSs. Our key insight involves using GPR to learn a DUR, using a user’s interaction history, that naturally embodies their diverse interest patterns. For any user $u$, let $\mathbf{V}_{u}=[\mathbf{v}_{u,1},\mathbf{v}_{u,2},...,\mathbf{v}_{u,l_{u}}]\in\mathbb{R}^{l_{u}\times d}$ be the embeddings of all items in their history. This is derived from their interaction list $\mathcal{V}_{u}$ and an item embedding matrix $\mathbf{V}$ (we describe how to obtain $\mathbf{V}$ in Sec. 4.4). Let $\mathbf{o}_{u}=[o_{u,v}|v\in\mathcal{V}_{u}]\in\mathbb{R}^{l_{u}}$ be the vector of $u$’s observed interactions with items in $\mathcal{V}_{u}$. We employ a GP to model $u$’s interests given the input points $\mathbf{V}_{u}$ and corresponding observations $\mathbf{o}_{u}$, $g_{u}\sim\mathcal{GP}(\mu_{u},k_{u})$, where, $g_{u}$, $\mu_{u}$ and $k_{u}$ are $u$’s personalized GP regressor, mean, and kernel function, respectively. The joint distribution of observation $\mathbf{o}_{u}$ and the predicted observation of a novel item $v_{*}$ is:
@@
$g_{u}(\mathbf{v})|\mathbf{o}_{u}\sim\mathcal{N}\big{(}\bar{g}_{u},\mathrm{cov}(g_{u})\big{)},$ (2)
@@
$\mathrm{cov}(g_{u})=$ $k(\mathbf{v},\!\mathbf{v})\!+\!\sigma_{*}^{2}\!\,-\!\mathbf{k}(\mathbf{v}_{*},\!\mathbf{V}_{u})[\mathbf{K}(\mathbf{V}_{u},\!\mathbf{V}_{u})\!+\!\sigma^{2}\!\mathbf{I}]^{-1}\mathbf{k}(\mathbf{V}_{u},\!\mathbf{v})^{T}\!.$ (4)
@@ 4.2 Retrieval List Generation
After obtaining a DUR $g_{u}$ for $u\in\mathcal{U}$ using GPR, we generate the retrieval list using the posterior $g_{u}(\mathbf{v})$ over all unobserved items. The top N items with the highest values form our retrieval list. We consider two methods for selection: (i) Thompson sampling (TS), a probabilistic method that selects items based on posterior sampling [15] and (ii) Upper Confidence Bound (UCB), a deterministic method that selects items based on their estimated rewards and uncertainties [14, 28]. The details for the retrieval list generation is shown in Appendix A.2.
@@ 4.3 GPR Parameter Tuning
The free parameters in our model are the kernel function $k$ and the standard deviation $\sigma$ in Eq. 1. We treat these as hyperparameters of GPR which are optimized using evaluation on a separate holdout set. Specifically, to generate $\mathcal{R}_{u}$ for user $u$, we fit the GPR model not to the complete interaction history $\mathcal{V}_{u}$, but to the reduced history $\mathcal{V}^{\text{h}}_{u}$, using the item embeddings and observed ratings. We assess retrieval performance using specific metrics (see Sec. 5.3) on both $\mathcal{R}_{u}$ and the holdout set $\mathcal{V}^{\text{d}}_{u}$ We tune GPR parameters using these criteria (we detail the adjustable parameters in Sec. 5.2).